A group of students got the following scores in a test: 6, 9, 12, 15, 18, and 21. Consider sample of size 4 that can be drawn from this population.


How many possible samples of size 4 can be drawn?


List all the possible samples and compute the sample mean.


Sample


Sample Mean


Construct the sampling distribution of the sample mean.


Sample Mean


Frequency


Probability P(x)


Draw a histogram corresponding to the sampling distribution of the sample mean.


1
Expert's answer
2021-05-07T14:15:51-0400

 The number of possible samples of size 4 can be calculated as: "_nC_r=\frac{n!}{r!\left(n-r\right)!}" where n = population size, r = sample size.

"_nC_r=\frac{720}{48}=15\:"

Number of possible samples of size 4 is 15.

To calculate the mean of the possible outcomes we use the formula:"Mean=\frac{sum\:of\:values\:of\:a\:possible\:otcome}{sample\:size}"



To calculate the Probability P (X) we use the formula: "Probability\:P\left(X\right)=\frac{Frequency}{Number\:of\:possible\:outcomes}"



A histogram corresponding to the sampling distribution of the sample mean.